# Beyond Classification: Towards Speech Emotion Reasoning with Multitask AudioLLMs

### v1

@@ Abstract
Audio Large Language Models (AudioLLMs) have achieved strong results in semantic tasks like speech recognition and translation, but remain limited in modeling paralinguistic cues such as emotion. Existing approaches often treat emotion understanding as a classification problem, offering little insight into the underlying rationale behind predictions. In this work, we explore emotion reasoning, a strategy that leverages the generative capabilities of AudioLLMs to enhance emotion recognition by producing semantically aligned, evidence-grounded explanations. To support this in multitask AudioLLMs, we introduce a unified framework combining reasoning-augmented data supervision, dual-encoder architecture, and task-alternating training. This approach enables AudioLLMs to effectively learn different tasks while incorporating emotional reasoning. Experiments on IEMOCAP and MELD show that our approach not only improves emotion prediction accuracy but also enhances the coherence and evidential grounding of the generated responses.
[…]
tcb@breakable
@@
footnotetext: Contributed equally to this work; authors are listed in alphabetical order.
@@ 2.1 AudioLLMs
Recent works explore enabling AudioLLMs to reason. Audio-CoT Ma et al. (2025) evaluates training-free chain-of-thought prompting, which requires AudioLLMs capable of general instruction following. Audio-Reasoner Xie et al. (2025) trains with a structured reasoning framework consisting of planning, captioning, reasoning and summarizing steps. Li et al. (2025) argues that the complex reasoning process in Audio-Reasoner may not be necessary, and the best practice remains an open research question.
@@ 2.2 Emotion recognition in AudioLLMs
In contrast to prior works, which focus primarily on improving emotion classification accuracy through architectural or input-level enhancements, our approach shifts the paradigm toward emotion reasoning. Rather than outputting a single emotion label, we leverage the generative capabilities of AudioLLMs to produce semantically coherent, evidence-grounded explanations for emotional states.
@@ 4.2 Datasets
For semantic tasks, we utilize the Spoken Question Answering (SQA) tasks in MNSC Wang et al. (2025), a corpus centered on Singlish, a Creole language rooted in English. We select MNSC because the pre-trained encoders and LLM are unlikely to have been exposed to its linguistic patterns during training, reducing the risk of performance bias from prior exposure. For further analysis in Section 6, we experiment with additional SQA tasks, such as Spoken-SQuAD Lee et al. (2018) and SLUE-P2-SQA5 Shon et al. (2023), and automatic speech recognition (ASR) tasks such as MNSC ASR Wang et al. (2025) and LibriSpeech Panayotov et al. (2015).
@@ 4.3 Evaluation
For emotion and sentiment recognition, we grade each emotion prediction on a binary scale, where a score of 1 indicates semantic alignment with the ground-truth label. Since AudioLLMs generate open-ended responses, traditional metrics such as exact match may be insufficient. The LLM-as-a-Judge approach allows us to assess the factual correctness and relevance of model outputs in a more flexible and human-aligned manner.
@@ 5.1 Effectiveness of reasoning-augmented training targets
[…]
@@
[…]
@@ 5.3 Quality of emotion reasoning responses
Table 4: Evaluation of model predictions with evidence-grounded reasoning predictions. Quotation Score measures the percentage of predictions containing at least one extractable quote. Groundedness and Relevance scores (0–100 scale) assess alignment with the transcript and support for the ground-truth emotion label, respectively.
@@ 5.4 Comparison with other models
We compare the emotion and sentiment recognition performance with end-to-end AudioLLMs evaluated in AudioBench: WavLLM Hu et al. (2023), Qwen2-Audio-7B-Instruct Chu et al. (2024), Phi-4-Multimodal-Instruct Abouelenin et al. (2025), MERaLiON-AudioLLM MERaLiON Team (2024), Qwen-Audio-Chat Chu et al. (2023), and SALMONN Tang et al. (2024). We also compare with cascaded models that process speech in sequential stages by converting audio to text using an automatic speech recognition module before feeding the transcript into a large language model: Whisper-Large-v2 with SEA-LIONv3 Ng et al. (2025), and Whisper-Large-v3 with Llama-3-8B-Instruct AI@Meta (2024). From Table 5, our proposed AudioLLM trained with reasoning-augmented targets achieves the best performance across IEMOCAP, MELD-ER and MELD-SR.
[…]
@@
We compare the performance of AudioLLM with emotion classifier Emotion2Vec+ Large in Table 6 on MELD. AudioLLM trained with reasoning-augmented targets achieves the highest weighted average accuracy (53.0%), outperforming both the label-only variant (47.1%) and Emotion2Vec+ Large (44.7%). However, the unweighted average remains comparable across models, suggesting that rare classes, such as disgust and fear, remain challenging.
It is important to note that comparisons with existing models are not controlled for differences in training data, which may influence performance outcomes.
Table 6: Class-wise emotion recognition performance comparison of Emotion2Vec+ Large classifier vs. AudioLLM (with Emotion2Vec+ Large emotion-centric encoder) trained on Label Only or Reasoning targets on MELD-ER. The unweighted average (Unwt Avg) treats all classes equally regardless of sample size, while the weighted average (Wt Avg) reflects the sample distribution by weighting each class accordingly.
@@ A.1 Datasets
For semantic tasks, we utilize the Multitask National Speech Corpus (MNSC) Wang et al. (2025), specifically SQA Part 3-6 and ASR Part 3-6, released under the Singapore Open Data License. We also use Spoken-SQuAD Lee et al. (2018) released under CC-BY-SA-4.0 License, SLUE-P2-SQA5 Shon et al. (2023) which is a collection of datasets released under CC-BY-SA-4.0 License and Apache License 2.0, and LibriSpeech Panayotov et al. (2015) released under CC-BY-4.0 License.
@@
A.2 AudioLLM responses
We conduct a qualitative analysis of model predictions across different reasoning formats on the IEMOCAP, MELD-ER and MELD-SR datasets. Each example includes the transcript, ground-truth label, and model-generated outputs under three supervision types: label-only, interpretive reasoning, and evidence-grounded reasoning. As shown in Table 8, 9 and 10, label-only responses often fail to capture the correct emotional nuance, defaulting to neutral predictions even when the emotion is apparent. In contrast, interpretive and evidence-grounded reasoning better align with the ground truth, offering richer justifications and improved emotion recognition. Notably, evidence-grounded reasoning demonstrates superior clarity by explicitly linking speech content and affective cues to the predicted emotion.
@@ A.2 AudioLLM responses
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ Abstract
Audio Large Language Models (AudioLLMs) have achieved strong results in semantic tasks like speech recognition and translation, but remain limited in modeling paralinguistic cues such as emotion. Existing approaches often treat emotion understanding as a classification problem, offering little insight into the underlying rationale behind predictions. In this work, we explore emotion reasoning, a strategy that leverages the generative capabilities of AudioLLMs to enhance emotion recognition by producing semantically aligned, evidence-grounded explanations. To support this in multitask AudioLLMs, we introduce a unified framework combining reasoning-augmented data supervision, dual-encoder architecture, and task-alternating training. This approach enables AudioLLMs to effectively learn different tasks while incorporating emotional reasoning. Experiments on IEMOCAP and MELD show that our approach not only improves emotion prediction accuracy but also enhances the coherence and evidential grounding of the generated responses. Experiments on two out-of-domain datasets demonstrate the generalization capabilities of the resulting model.
@@ 2.1 AudioLLMs
Recent works explore enabling AudioLLMs to reason. Audio-CoT Ma et al. (2025) evaluates training-free chain-of-thought prompting, which requires AudioLLMs capable of general instruction following. Audio-Reasoner Xie et al. (2025) trains with a structured reasoning framework consisting of planning, captioning, reasoning and summarizing steps. Li et al. (2025a) argues that the complex reasoning process in Audio-Reasoner may not be necessary, and the best practice remains an open research question.
@@ 2.2 Emotion recognition in AudioLLMs
In contrast to prior works, which focus primarily on improving emotion classification accuracy through architectural or input-level enhancements, our approach shifts the paradigm toward emotion reasoning. Rather than outputting a single emotion label, we leverage the generative capabilities of AudioLLMs to produce semantically coherent, evidence-grounded explanations.
@@ 4.2 Datasets
For semantic tasks, we utilize the Spoken Question Answering (SQA) tasks in MNSC Wang et al. (2025a), a corpus centered on Singlish, a Creole language rooted in English. We select MNSC because the pre-trained encoders and LLM are unlikely to have been exposed to its linguistic patterns during training, reducing the risk of performance bias from prior exposure. For further analysis in Section 6, we experiment with additional SQA tasks, such as Spoken-SQuAD Lee et al. (2018) and SLUE-P2-SQA5 Shon et al. (2023), and automatic speech recognition (ASR) tasks such as MNSC ASR Wang et al. (2025a) and LibriSpeech Panayotov et al. (2015).
@@ 4.3 Evaluation
For emotion and sentiment recognition, we grade each emotion prediction on a binary scale, where a score of 1 indicates semantic alignment with the ground-truth label. Since AudioLLMs generate open-ended responses, traditional metrics such as exact match may be insufficient. The LLM-as-a-Judge approach allows us to assess the factual correctness and relevance of model outputs in a more flexible manner. We analyze the evaluation approach in Section  7.
@@ 5.1 Effectiveness of reasoning-augmented training targets
Concat Dim Training ER/SR SQA IEMOCAP MELD-ER MELD-SR Avg Part 3 Part 4 Part 5 Part 6 Avg Overall Avg None Joint 43.1 51.9 61.9 52.3 49.4 48.4 57.6 62.4 54.5 53.5 Feature Joint 40.4 51.5 62.6 51.5 47.2 50.6 58.6 63.8 55.1 53.5 Alt 1 epoch 53.7 54.6 62.0 56.8 49.2 50.0 56.4 63.0 54.7 55.6 Alt 4 epochs 48.5 52.8 61.4 54.2 42.4 40.4 57.0 62.4 50.6 52.1 Sequence Joint 44.1 48.8 58.1 50.3 32.8 30.8 40.2 47.6 37.9 43.2 Alt 1 epoch 56.6 52.4 60.7 56.6 48.0 50.6 59.2 63.4 55.3 55.8 Alt 4 epochs 55.1 52.0 61.5 56.2 52.4 49.4 57.6 64.4 56.0 56.1
@@
Training PL Encoder ER/SR SQA IEMOCAP MELD-ER MELD-SR Avg Part 3 Part 4 Part 5 Part 6 Avg Overall Avg Alt 1 epoch Whisper-Large (637M) 56.6 52.4 60.7 56.6 48.0 50.6 59.2 63.4 55.3 55.8 Whisper-Small (88M) 56.8 52.5 60.6 56.6 50.8 50.4 58.6 61.4 55.3 55.9 Whisper-Tiny (8M) 46.2 49.3 53.1 49.5 25.8 23.8 27.0 32.0 27.2 36.7 HuBERT-XL (962M) 51.3 46.1 51.9 49.8 26.8 23.6 27.0 28.6 26.5 36.5 Emotion2Vec+ Large (164M) 57.1 50.0 56.6 54.6 28.2 22.4 28.2 32.3 27.8 39.3 Emotion2Vec+ base (93M) 63.5 45.9 55.0 54.8 24.6 24.2 29.4 29.6 27.0 38.9 Alt 4 epochs Whisper-Large (637M) 55.1 52.0 61.5 56.2 52.4 49.4 57.6 64.4 56.0 56.1 Whisper-Small (88M) 55.3 52.6 61.6 56.5 48.6 49.2 61.4 61.8 55.3 55.8 Whisper-Tiny (8M) 50.3 52.7 59.7 54.2 47.8 47.0 58.0 62.2 53.8 54.0 HuBERT-XL (962M) 44.0 49.3 59.0 50.8 49.6 47.8 59.0 62.2 54.7 53.0 Emotion2Vec+ Large (164M) 63.8 53.0 61.1 59.3 49.2 51.2 59.0 62.8 55.6 57.2 Emotion2Vec+ base (93M) 56.3 52.0 60.4 56.2 46.2 45.8 54.4 61.0 51.9 53.7
@@ 5.3 Quality of emotion reasoning responses
Table 4: Evaluation of evidence-grounded reasoning. Quotation Score measures the percentage of predictions containing at least one extractable quote. Groundedness and Relevance scores (0–100 scale) assess alignment with the transcript and support for the ground-truth emotion label, respectively.
@@ 5.4 Comparison with other models
We compare the emotion and sentiment recognition performance with end-to-end AudioLLMs evaluated in AudioBench: WavLLM Hu et al. (2023), Qwen2-Audio-7B-Instruct Chu et al. (2024), Phi-4-Multimodal-Instruct Abouelenin et al. (2025), MERaLiON-AudioLLM MERaLiON Team (2024), Qwen-Audio-Chat Chu et al. (2023), SALMONN Tang et al. (2024), R1-AQA Li et al. (2025b) and Audio-Reasoner Xie et al. (2025). R1-AQA is trained with reinforcement learning for improved thinking capabilities, and Audio-Reasoner is trained for planning and reasoning. We also compare with cascaded models that process speech in sequential stages by converting audio to text using an automatic speech recognition module before feeding the transcript into a large language model: Whisper-Large-v2 with SEA-LIONv3 Ng et al. (2025), and Whisper-Large-v3 with Llama-3-8B-Instruct AI@Meta (2024). From Table 5, our proposed AudioLLM-Reasoning achieves the best performance for IEMOCAP, MELD-SR and overall.
Model IEMOCAP MELD-ER MELD-SR Avg Audio-Reasoner 51.0 55.9 54.5 53.8 WavLLM 59.8 41.6 51.1 50.8 Qwen2-Audio 54.0 41.6 53.9 49.8 Cascade: Whisper+SEA-LION 44.3 47.4 56.6 49.4 R1-AQA 57.2 42.8 40.7 46.9 Phi-4-Multimodal 41.0 43.5 51.6 45.4 MERaLiON 48.5 36.4 46.2 43.7 Cascade: Whisper+Llama3 46.7 36.8 45.6 43.0 Qwen-Audio 29.4 50.7 44.9 41.7 SALMONN 23.8 30.5 41.8 32.0 AudioLLM-Reasoning 63.8 53.0 61.1 59.3
@@
5.5 Generalization to out-of-domain datasets
We evaluate the generalizability of the proposed model on two out-of-domain emotion recognition datasets not seen at training: M3ED Zhao et al. (2022) and CPQA Wang et al. (2025b). M3ED contains utterances from Mandarin Chinese TV series. CPQA is a contexual paralinguistic question-answering dataset constructed using speech data collected from top Singapore YouTube channels; we use only the emotion recognition set for our evaluation. Both datasets are annotated for seven emotion classes, namely neutral, happy, disgust, sad, surprise, anger and fear. We compare our model with emotion classifier Emotion2Vec+ Large, R1-AQA which is trained with reinforcement learning for improved thinking capabilities, and Audio-Reasoner which is trained for planning and reasoning. From Table 6, our proposed AudioLLM-Reasoning achieves the best performance on both M3ED and CPQA-ER. Additional class-wise comparisons with Emotion2Vec+ Large are in Appendix A.2.
Model M3ED CPQA-ER Avg Audio-Reasoner 45.2 48.5 46.9 Emotion2Vec+ Large 47.9 37.9 42.9 R1-AQA 38.4 43.1 40.8 AudioLLM-Reasoning 48.6 49.0 48.8
Table 6: Emotion recognition performance on out-of-domain datasets.
@@ 6 Further Analysis
7 Assessment of LLM-as-a-Judge metric
We adopt the LLM-as-a-Judge metric for evaluation as the AudioLLMs can generate open-ended and expressive outputs. Traditional metrics relying on exact string matching are insufficient in this context. We identify three main cases where traditional metrics can fail to capture the true quality of model responses: 1. the model uses semantically equivalent but lexically different expressions to describe emotions; 2. the model output includes multiple plausible or related emotional states; and 3. the predicted labels fall into semantically overlapping categories (e.g., excited vs. happy, or anger vs. frustration). Outside of these special cases, the LLM-as-a-Judge metric effectively reduces to an accuracy-like measure, where the model’s prediction is compared against a reference by string matching.
Out of the 858 test samples in IEMOCAP, 27.6% of the AudioLLM-Reasoning emotion predictions fall into the special cases. Out of the 2610 test samples in MELD-ER, 6% of the AudioLLM-Reasoning emotion predictions fall into the special cases. To better understand the characteristics of the LLM-as-a-Judge metric, we randomly selected 50 special-case samples each from the IEMOCAP and MELD-ER datasets to be assessed by 4 human evaluators. The evaluators are provided with the audio clips, references and model’s answers, and are instructed to rate the model’s answers based on their alignment with the audio clips and references. For both IEMOCAP and MELD-ER, only 2% of the selected samples are scored as correct by the LLM judge but scored as wrong by the human evaluators. In contrast, 16% of the IEMOCAP selected samples and 30% of the MELD-ER selected samples are scored as wrong by the LLM judge but scored as correct by the human evaluators, suggesting that the LLM judge tends to be more conservative or strict in its assessments compared to human evaluators.
@@ A.1 Datasets
For semantic tasks, we utilize the Multitask National Speech Corpus (MNSC) Wang et al. (2025a), specifically SQA Part 3-6 and ASR Part 3-6, released under the Singapore Open Data License. We also use Spoken-SQuAD Lee et al. (2018) released under CC-BY-SA-4.0 License, SLUE-P2-SQA5 Shon et al. (2023) which is a collection of datasets released under CC-BY-SA-4.0 License and Apache License 2.0, and LibriSpeech Panayotov et al. (2015) released under CC-BY-4.0 License.
@@
A.2 Comparison with Emotion2Vec+
We compare the performance of AudioLLM with emotion classifier Emotion2Vec+ Large. For the in-domain dataset MELD-ER in Table 8, AudioLLM trained with reasoning-augmented targets achieves the highest weighted average score (53.0%), outperforming both the label-only variant (47.1%) and Emotion2Vec+ Large (44.7%). As the AudioLLMs are trained on a limited set of emotion datasets (i.e. IEMOCAP and MELD), they tend to fit to the training label distribution. In contrast, Emotion2Vec+ Large is trained on five emotion datasets.
The performance on the out-of-domain dataset CPQA-ER in Table 9 is affected less by the training label distribution. AudioLLM-Reasoning outperforms Emotion2Vec+ Large classifier on 6 out of 7 classes, and significantly outperforms the classifer on both the unweighted average score (45.3% vs. 36.0%) and the weighted average score (49.7% vs. 38.3%). Moreover, the discriminative classifier cannot readily extend beyond its label space. In Table 9, 11 samples involving states such as frustration, embarrassment, and mixture of emotions are excluded. The vanilla AudioLLM trained with label-only targets lacks generalization capabilities and has severely degraded performances.
Table 8: In-domain comparison: Class-wise emotion recognition performance comparison of Emotion2Vec+ Large classifier vs. AudioLLM (with Emotion2Vec+ Large emotion-centric encoder) trained on Label Only or Reasoning targets on MELD-ER. The unweighted average (Unwt Avg) treats all classes equally regardless of sample size, while the weighted average (Wt Avg) reflects the sample distribution by weighting each class accordingly.
Class Num Samples Emotion2Vec+ Large AudioLLM - Label Only AudioLLM - Reasoning Neutral 102 12.7 13.7 32.4 Joy 91 71.4 38.5 74.7 Disgust 30 13.3 10.0 20.0 Sadness 57 59.6 22.8 40.4 Surprise 54 31.5 1.9 33.3 Anger 83 43.4 28.9 78.3 Fear 50 20.0 4.0 38.0 Unwt Avg 467 36.0 17.1 45.3 Wt Avg 467 38.3 19.7 49.7
Table 9: Out-of-domain comparison: Class-wise emotion recognition performance comparison of Emotion2Vec+ Large classifier vs. AudioLLM (with Emotion2Vec+ Large emotion-centric encoder) trained on Label Only or Reasoning targets on CPQA-ER. 11 samples with ground-truth labels outside of the label space of Emotion2Vec+ Large classifier have been excluded.
A.3 AudioLLM responses
We conduct a qualitative analysis of model predictions across different reasoning formats on the IEMOCAP, MELD-ER and MELD-SR datasets. Each example includes the transcript, ground-truth label, and model-generated outputs under three supervision types: label-only, interpretive reasoning, and evidence-grounded reasoning. As shown in Table 10, 11 and 12, label-only responses often fail to capture the correct emotional nuance, defaulting to neutral predictions even when the emotion is apparent. In contrast, interpretive and evidence-grounded reasoning better align with the ground truth, offering richer justifications and improved emotion recognition. Notably, evidence-grounded reasoning demonstrates superior clarity by explicitly linking speech content and affective cues to the predicted emotion.
@@ A.3 AudioLLM responses
Label Only: The speaker’s speech reflects a neutral state
@@
Label Only: The speaker’s speech suggesting a neutral state.
@@
Label Only: The speaker’s speech suggesting a neutral state.
@@
Label Only: The speaker’s emotion seems neutral.
@@
Label Only: The speaker’s emotion seems neutral.
@@
Label Only: The speaker’s speech reflects neutral.
@@
Label Only: The speaker’s sentiment is neutral, based on their speech.
@@
Label Only: The speaker’s sentiment is neutral, based on their speech.